# Bayesian Transfer Operators in Reproducing Kernel Hilbert Spaces

### v1

@@ 1 Introduction
Transfer operator theory and reproducing kernel Hilbert spaces (RKHSs) are popular approaches that result in analytically informative algorithms (Mauroy et al., 2020; Klus et al., 2016; Brunton et al., 2022; Rasmussen and Williams, 2006; Shawe-Taylor, 2004; Berlinet and Thomas-Agnan, 2011; Cressie, 1990; Hofmann et al., 2008; Schölkopf and Smola, 2002), while also contributing widely to practical applications (Budišić et al., 2012; Saitoh et al., 2016; Rowley et al., 2009; Susuki and Mezic, 2011; Susuki et al., 2011; Berger et al., 2015; Bruder et al., 2019; Netto and Mili, 2018). Both Koopman operator and RKHS techniques frame their corresponding problems in a functional analytic setting. Therefore, it stands to reason that by combining methods from both of these two fields, we can design robust algorithms that provide compact and interpretable models for potentially stochastic time-series.
@@ 1.1 Dynamic Mode Decomposition
The Koopman operator, being an infinite-dimensional object, must often be empirically approximated before it can be implemented on a computer. This involves discretizing the operator and identifying a suitable set of basis functions that span a finite-dimensional subspace wherein the spectral properties of the operator can be represented – a task that falls under the heading of dictionary learning (Li et al., 2017; Takeishi et al., 2017; Wehmeyer and Noé, 2018; Yeung et al., 2019; Azencot et al., 2020; Eivazi et al., 2021; Mardt et al., 2018).
@@
In part, DMD remains attractive because it is based on an eigenvalue decomposition of a model obtained through linear regression. After projecting the system to a higher-dimensional space, DMD can provide accurate models of nonlinear (Williams et al., 2015a, b) and stochastic systems (Klus et al., 2020a; Wanner and Mezic, 2022; Črnjarić-Žic et al., 2020; Colbrook et al., 2024). In other words, all the techniques belonging to linear time-invariant ODEs, such as spectral analysis, multi-step forecasting, numerical simulation, and even control (Peitz and Klus, 2019; Peitz et al., 2020; Proctor et al., 2016; Otto and Rowley, 2021) can be readily transferred to nonlinear systems. Moreover, should the system contain a large number of variables (e.g., in fluid dynamics), DMD can facilitate dimensionality reduction.
@@ 1.2 Gaussian Process Regression
Following the authors Williams et al. (2015b); Klus et al. (2018); DeGennaro and Urban (2019); Fujii and Kawahara (2019); Das and Giannakis (2020); Das et al. (2021); Ikeda et al. (2022); Bevanda et al. (2024); Klus et al. (2020b); Baddoo et al. (2022); Bevanda et al. (2025), the underlying machinery of our DMD algorithm is a reproducing kernel function (i.e., a covariance function). Kernel methods are non-parametric, enhance interpretability of features, and are generally more flexible compared to parametric models (e.g., multilayer perceptrons), since they encompass all functions sharing the same degree of smoothness. It is the non-parametric nature of kernel algorithms that grant them the intrinsic ability to implicitly construct a functional basis for a model directly from data. Thereby, they provide a natural choice for the hypotheses space: an RKHS. It is these properties that often lead to strong theoretical performance guarantees of convergence and statistical consistency (Schölkopf and Smola, 2002; Klus et al., 2020c; Philipp et al., 2024).
One popular kernel regression framework (for modeling dynamical systems), is Gaussian process (GP) regression (MacKay et al., 1998; Rasmussen and Williams, 2006; Girard et al., 2002; Wang et al., 2005; Solak et al., 2002). GP regression is a function approximation technique that is capable of providing accurate estimations of an unknown function based on a limited set of noisy training targets and high-level assumptions about the function. Moreover, a GP model is numerically straightforward to estimate, the predictions of the model are fully probabilistic, and the theory provides a well-founded framework for model selection.
@@ 1.3 Contributions of this work
The idea is to transfer the attractive features of GP regression to DMD. Thereby, we develop a Bayesian interpretation of the embedded Perron–Frobenius operator (Klus et al., 2020b) which opens up avenues to extend kernel-based DMD algorithms with established sparsity and hyperparameter tuning techniques. The result is a modification of ‘Extended DMD’ (EDMD; Williams et al. (2015a)) that constructs models from noisy measurement data of a nonlinear dynamical system and provides uncertainty estimates for multi-step predictions.
The remainder of this paper is structured as follows. Section 2 introduces a Bayesian formulation of DMD, and Section 3 explores its mathematical consequences and numerical performance. Lastly, in Section 4, we conclude with some final remarks, open questions, and point toward straightforward extensions. In Appendix A we review (sparse) GP regression and the notation of Rasmussen and Williams (2006), and for convenience and completeness we also present an overview of (kernel) transfer operators (Klus et al., 2020b) in Appendix B.
@@ 2 Gaussian Process Dynamic Mode Decomposition
Because kernel-based DMD algorithms often quickly become impractical for large data sets due to the high training cost, one of our goals is to apply the variational free energy (VFE) method (Titsias, 2009) to kernel Koopman operators (Klus et al., 2020b). For example, in Klus et al. (2020b), to estimate the Koopman matrix we need to take the inverse of a Gramian which has a complexity of $\mathcal{O}(N^{3})$; moreover, the eigenvalue decomposition also has a complexity of $\mathcal{O}(N^{3})$. Therefore, we are motivated to find a better trade-off between the model’s generalization error and its computational complexity. We want a non-parametric model that automatically adapts to the complexity of the dynamical system, but not to the point where all the training inputs define the basis dictionary.
@@
For example, EDMD does not intrinsically provide an optimization scheme for selecting either the hyperparameters or the dictionary, and therefore requires another explicit strategy (Tabish et al., 2025). Neither does EDMD provide uncertainty-aware predictions, nor does it take into account measurement noise from the outset. Our amalgamation of DMD and GP regression wraps up the entire model inference pipeline into one coherent theory that simultaneously addresses hyperparameter optimization and sparse dictionary learning. This naturally raises the question of whether such an approach is overly restrictive and sacrifices the flexibility of frequentist methods.
@@ 2.1 Operators as Random Variables
Pseudo-inputs
$\mathbf{Z}\coloneqq[\bm{z}_{1},...,\bm{z}_{M}]$
Snapshot matrices
$\mathbf{X}\coloneqq[\bm{x}_{1},...,\bm{x}_{N}]$
@@
Dictionaries
$\Phi_{Z}\coloneqq[\varphi(\bm{z}_{1}),...,\varphi(\bm{z}_{M})]$
@@
Covariance functions
$\kappa_{\mathrm{pr}}(\bm{x},\bm{x}^{\prime})\coloneqq\langle\varphi(\bm{x}),\varphi(\bm{x}^{\prime})\rangle_{\mathbb{H}}$
@@
Covariance matrices
$\mathbf{K}_{ZZ}\coloneqq\kappa_{\mathrm{pr}}(\mathbf{Z},\mathbf{Z})$
@@
Covariance operators
$\Xi_{XX}\coloneqq\kappa_{\mathrm{pr}}(\mathbf{X},\mathbf{X})\otimes\mathcal{C}_{\textrm{bc}}$
@@
Feature vectors
$\mathbf{k}_{Z}(\bm{x})\coloneqq\Phi_{Z}^{\top}\varphi(\bm{x})$
@@
Gramian matrix
$\mathbf{C}_{XX}\coloneqq\mathbf{K}_{ZX}\,\mathbf{K}_{ZX}^{\top}$
Stiffness matrix
$\mathbf{C}_{XY}\coloneqq\mathbf{K}_{ZX}\,\mathbf{K}_{ZY}^{\top}$
Regularization
$\tilde{\mathbf{K}}_{XX}\coloneqq{\mathbf{K}}_{XX}+\sigma_{Y}^{2}\mathbf{I}_{N}$
@@
We start with a full Bayesian treatment of the embedded Perron–Frobenius operator (PFO) (Klus et al., 2020b), which we denote by $\mathcal{P}_{\varepsilon}:\mathbb{H}\rightarrow\mathbb{H}$. The embedded Perron–Frobenius operator, is the adjoint of the kernel Koopman operator. Unlike the PFO which propagates densities, the embedded PFO can be thought of as a push-forward map that acts on representations of probability densities. In short, the idea of feature maps is extended to the space of probability distributions by thinking of $\varphi(\bm{x})\,\forall\,\bm{x}\in\mathbb{X}$ as an injective representation of a density function. See Appendix C for a more detailed discussion of kernel mean embeddings.
Let us consider the ‘implicitly’ lifted training data set obtained from a dynamical system as $\left(\Phi_{X},\Phi_{Y}\right)=\{\left(\varphi\left(\bm{x}_{i}\right),\varphi\left(\bm{y}_{i}\right)\right)\}_{i=1}^{N}$, where the targets $\{\bm{y}_{i}\}_{i=1}^{N}$ are produced by the discrete-time flow map $\bm{F}:\mathbb{X}\rightarrow\mathbb{X}$, and $\mathbb{X}\subseteq\mathbb{R}^{D}$. We further assume only isotropic measurement noise on the targets:
@@
For the sake of clarity, allow a brief oversimplification and consider the D=1D\!=\!1 case. In this setting, the feature map $\varphi(\bm{x})\!\coloneqq\kappa_{\textrm{pr}}(\bm{x},\cdot)\!\in\!\mathbb{H}$ is defined by the GP prior over a one-dimensional flow map, i.e. a real- and scalar-valued kernel $\kappa_{\textrm{pr}}$.
@@
Lifted Observation Model. We assume that a noisy lifted target $\varphi(\bm{y})\in\mathbb{H}$ can be approximated by the sum of the latent noise-free value $\mu_{Y\mid\bm{x}}\coloneqq\mathcal{P}_{\varepsilon}\;\varphi(\bm{x})\in\mathbb{H}$ and zero mean Gaussian process noise:
@@
where $\nu_{i}\sim\mathcal{GP}\left(0,\sigma_{Y}^{2}\kappa_{\textrm{bc}}\right)$ models noise in the (infinite-dimensional) RKHS $\mathbb{H}$.
Definition 2.1 assumes that the lifted targets are also modeled as scalar-valued GPs, and therefore the likelihood can be expressed as $p(\varphi(\bm{y})\mid\mathcal{P}_{\varepsilon},\bm{x})=\mathcal{GP}(\mu_{Y\mid\bm{x}},\sigma_{Y}^{2}\kappa_{\textrm{bc}})$.
The purpose of the Bayesian-consistency kernel $\kappa_{\textrm{bc}}$ is to relate the observation model in the state space to the observation model in the feature space. It models the correlations between the ‘components’ of $\varphi(\cdot)$ such that consistency is maintained with the flow map. The associated covariance operator $\mathcal{C}_{\textrm{bc}}:\mathbb{H}\rightarrow\mathbb{H}$, which in practice is required to be self-adjoint and invertible, may be thought of as representing the statistical properties of the full-state observables within the RKHS. In Section 2.3 we will discretize this operator.
Similar to scalar-valued GP regression, our prior beliefs about the statistical properties of the operator $\mu_{Y\mid\cdot}:\mathbb{X}\rightarrow\mathbb{H}$ are represented by a GP, $\mu_{Y\mid\cdot}\sim\mathcal{GP}^{\infty}\left(\mathcal{M}_{\textrm{pr}},{\xi}_{\textrm{pr}}\right)$. However, here we are defining the prior covariance ${\xi}_{\textrm{pr}}\colon\mathbb{X}\times\mathbb{X}\to\mathbb{H}\,\otimes\,\mathbb{H}$ as an operator-valued kernel. Then, without loss of generality, we choose the mean operator $\mathcal{M}_{\textrm{pr}}:\mathbb{X}\rightarrow\mathbb{H}$ such that it produces a function that is always identically zero. It follows that the prior distribution over the latent operator evaluations $\mu_{Y\mid\mathbf{X}}\coloneqq\mathcal{P}_{\varepsilon}\;\Phi_{X}\in\mathbb{H}^{1\times N}$ is
@@
where $\Xi_{XX}$ is an element of $(\mathbb{H}\,\otimes\,\mathbb{H})^{N\times N}$.
The Riesz–Fréchet representation theorem (Hsing and Eubank, 2015; Fréchet, 1904) says that for each $\bm{x}\!\in\!\mathbb{X}$ and $g\!\in\!\mathbb{H}$, there is an element $\Xi(\bm{x})g\!:\mathbb{X}\rightarrow\mathbb{H}$, such that the reproducing property is
@@
This implies that the feature map $\Xi(\bm{x})$ is an element of the space of bounded linear operators $\mathfrak{B}(\mathbb{H},\mathbb{H}^{\mathbb{X}})$, and $\mathbb{H}^{\mathbb{X}}=\mathfrak{B}(\mathbb{X},\mathbb{H})$ is the feature space. Thereby, the relationship between the reproducing kernel $\xi_{\textrm{pr}}(\cdot,\cdot)$ and its associated feature map $\Xi(\bm{x})$ is
@@
The entry at $\left(\xi_{\textrm{pr}}\left(\bm{x},\bm{x}^{\prime}\right)\right)_{\bm{y},\bm{y}^{\prime}}$ determines the prior covariance between $\mu_{Y\mid\bm{x}}(\bm{y})$ and $\mu_{Y\mid\bm{x}^{\prime}}(\bm{y}^{\prime})$, i.e., it expresses the degree to which the response at $(\bm{x},\bm{y})$ is affected by $(\bm{x}^{\prime},\bm{y}^{\prime})$.
Given the GP likelihood, the posterior retains the same structure as in the scalar-valued case (Alvarez et al., 2012; Micchelli and Pontil, 2004, 2005):
$\mathcal{M}_{\textrm{pst}}(\bm{x})$ $\coloneqq\xi_{X}^{\top}({\bm{x}})\text{ }\mathcal{A}=\sum_{i=1}^{N}{\xi}_{\textrm{pr}}(\bm{x}_{i},\bm{x})\mathcal{A}_{i},$ (2.1)
@@
We unpack these equations by concatenating the noisy targets ‘end-to-end’ $\Upsilon^{\top}\!\coloneqq\!\operatorname{vec}(\Phi_{Y})\!\in\!\mathbb{H}^{\oplus N}$. We denote the weight operator by $\mathcal{A}\coloneqq\tilde{\Xi}_{XX}^{-1}\text{ }\Upsilon^{\top}$, the noisy covariance operator by $\tilde{\Xi}_{XX}\coloneqq\Xi_{XX}+\Sigma$, and the feature “vector" by $\mathbf{\xi}_{X}(\bm{x})\coloneqq[{\xi}_{\textrm{pr}}(\bm{x}_{1},\bm{x}),\ldots,{\xi}_{\textrm{pr}}(\bm{x}_{N},\bm{x})]^{\top}$.
While perhaps at present it isn’t precisely clear why, the structure of the lifted model restricts the operator-valued kernel to $\xi_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\!=\!\kappa_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\,\mathcal{C}_{\textrm{bc}}$. This form allows us to express the covariance operators in terms of the prior covariances over latent values of the flow map and Kronecker tensor products with $\mathcal{C}_{\textrm{bc}}$, i.e., $\Xi_{XX}\coloneqq\mathbf{K}_{XX}\otimes\mathcal{C}_{\textrm{bc}}$, $\Xi_{ZX}\coloneqq\mathbf{K}_{ZX}\otimes\mathcal{C}_{\textrm{bc}}$. Similarly, the “measurement" noise covariance operator is denoted by $\Sigma\coloneqq\mathbf{I}_{N}\,\otimes\,\sigma_{Y}^{2}\,\mathcal{C}_{\textrm{bc}}$. Thereby, we are in essence implementing an intrinsic model of coregionalization (Alvarez et al., 2012).
@@
Hence, we recover the same expression for ${\mathcal{P}}_{\varepsilon}$ derived in (Klus et al., 2020b), while also mirroring the frequentist approach of (Grünewälder et al., 2012). In the expression of the mean, the operator $\mathcal{C}_{\textrm{bc}}$ plays no role whatsoever, yet we observe that the posterior covariance over ${\mathcal{P}}_{\varepsilon}\varphi(\bm{x})$ retains the same form as the prior: $\xi_{\textrm{pst}}(\bm{x},\bm{x}^{\prime})=\kappa_{\textrm{pst}}(\bm{x},\bm{x}^{\prime})\,\mathcal{C}_{\textrm{bc}}$, i.e., a linear transformation of the covariance over the flow map.
In the case where $D>1$, we may want to treat the prior kernel $\kappa_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})$ as matrix-valued and include anisotropic measurement noise (Alvarez et al., 2012; Wilson et al., 2012). However, for the rest of this manuscript we will continue assuming that the channels in the flow map are unrelated and that the sensor noise is isotropic.

### v2

@@ 1 Introduction
Transfer operator theory and reproducing kernel Hilbert spaces (RKHSs) are popular approaches that result in analytically informative algorithms (Mauroy et al., 2020; Klus et al., 2016; Brunton et al., 2022; Rasmussen and Williams, 2006; Shawe-Taylor, 2004; Berlinet and Thomas-Agnan, 2011; Cressie, 1990; Hofmann et al., 2008; Schölkopf and Smola, 2002), while also contributing widely to practical applications (Budišić et al., 2012; Saitoh et al., 2016; Rowley et al., 2009; Susuki and Mezic, 2011; Susuki et al., 2011; Berger et al., 2015; Bruder et al., 2019; Netto and Mili, 2018). Both Koopman operator and RKHS techniques frame their corresponding problems in a functional analytic setting. Therefore, it stands to reason that by combining methods from both of these two fields, we can design robust algorithms that provide compact and interpretable models for (stochastic) time-series.
@@ 1.1 Dynamic Mode Decomposition
The Koopman operator, being an infinite-dimensional object, must often be empirically approximated before it can be implemented on a computer. This involves discretizing the operator and identifying a suitable set of basis functions that span a finite-dimensional subspace wherein the spectral properties of the operator can be represented — a task that falls under the heading of dictionary learning (Li et al., 2017; Takeishi et al., 2017; Wehmeyer and Noé, 2018; Yeung et al., 2019; Azencot et al., 2020; Eivazi et al., 2021; Mardt et al., 2018).
@@
In part, DMD remains attractive because it is based on an eigenvalue decomposition of a model obtained through linear regression. After projecting the system to a higher-dimensional space, DMD can provide accurate models of nonlinear (Williams et al., 2015a, b) and stochastic systems (Klus et al., 2020b; Wanner and Mezic, 2022; Črnjarić-Žic et al., 2020; Colbrook et al., 2024). In other words, all the techniques belonging to linear time-invariant ODEs, such as spectral analysis, multi-step forecasting, numerical simulation, and even control (Peitz and Klus, 2019; Peitz et al., 2020; Proctor et al., 2016; Otto and Rowley, 2021) can be readily transferred to nonlinear systems. Moreover, should the system contain a large number of variables (e.g., in fluid dynamics), DMD can facilitate dimensionality reduction.
@@ 1.2 Gaussian Process Regression
Following the authors Williams et al. (2015b); Klus et al. (2018); DeGennaro and Urban (2019); Fujii and Kawahara (2019); Das and Giannakis (2020); Das et al. (2021); Ikeda et al. (2022); Bevanda et al. (2024); Klus et al. (2020c); Baddoo et al. (2022); Bevanda et al. (2025), the underlying machinery of our DMD algorithm is a reproducing kernel function (i.e., a covariance function). Kernel methods are non-parametric, enhance interpretability of features, and are generally more flexible compared to parametric models (e.g., multilayer perceptrons), since they encompass all functions sharing the same degree of smoothness. It is the non-parametric nature of kernel algorithms that grant them the intrinsic ability to implicitly construct a functional basis for a model directly from data. Thereby, they provide a natural choice for the hypotheses space: an RKHS. It is these properties that often lead to strong theoretical performance guarantees (Schölkopf and Smola, 2002; Klus et al., 2020a; Philipp et al., 2024).
One popular kernel regression framework (for modeling dynamical systems), is Gaussian process (GP) regression (MacKay and others, 1998; Rasmussen and Williams, 2006; Girard et al., 2002a; Wang et al., 2005; Solak et al., 2002). GP regression is a function approximation technique that is capable of providing accurate estimations of an unknown function based on a limited set of noisy training targets and high-level assumptions about the function. Moreover, a GP model is numerically straightforward to estimate, the predictions of the model are fully probabilistic, and the theory provides a well-founded framework for model selection.
@@ 1.3 Contributions of this work
The core idea is to transfer the attractive features of GP regression to DMD. Thereby, we develop a Bayesian interpretation of the embedded Perron–Frobenius operator (Klus et al., 2020c) which opens up avenues to extend kernel-based DMD algorithms with established sparsity and hyperparameter tuning techniques. The result is a modification of ‘Extended DMD’ (EDMD; Williams et al. (2015a)) that constructs models from noisy measurement data of a nonlinear dynamical system and provides uncertainty estimates for multi-step predictions.
The remainder of this paper is structured as follows. Section 2 introduces a Bayesian formulation of DMD, and Section 3 explores its mathematical consequences and numerical performance. Lastly, in Section 4, we conclude with some final remarks, open questions, and point toward straightforward extensions. In Appendix A we review (sparse) GP regression and the notation of Rasmussen and Williams (2006), and for convenience and completeness we also present an overview of (kernel) transfer operators (Klus et al., 2020c) in Appendix B.
@@ 2 Gaussian Process Dynamic Mode Decomposition
Because kernel-based DMD algorithms often quickly become impractical for large data sets due to the high training cost, one of our goals is to apply the variational free energy (VFE) method (Titsias, 2009) to kernel Koopman operators (Klus et al., 2020c). For example, in Klus et al. (2020c), to estimate the Koopman matrix we need to take the inverse of a Gramian which has a complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of training samples; moreover, the eigenvalue decomposition also has a complexity of $\mathcal{O}(N^{3})$. Therefore, we are motivated to find a better trade-off between the model’s generalization error and its computational complexity. We want a non-parametric model that automatically adapts to the complexity of the dynamics, but not to the point where all the training inputs define the basis dictionary.
@@
For example, EDMD does not intrinsically provide an optimization scheme for selecting either the hyperparameters or the dictionary, and therefore requires another explicit strategy, e.g., (Tabish et al., 2025). Neither does EDMD provide uncertainty-aware predictions, nor does it take into account measurement noise from the outset. Our amalgamation of DMD and GP regression wraps up the entire model inference pipeline into one coherent theory that simultaneously addresses hyperparameter optimization and sparse dictionary learning. This naturally raises the question of whether such an approach is overly restrictive and sacrifices the flexibility of frequentist methods.
@@ 2.1 Operators as Random Variables
Pseudo-inputs $\mathbf{Z}\coloneqq[\bm{z}_{1},...,\bm{z}_{M}]$
Snapshot matrices $\mathbf{X}\coloneqq[\bm{x}_{1},...,\bm{x}_{N}]$
@@
Dictionaries $\Phi_{Z}\coloneqq[\varphi(\bm{z}_{1}),...,\varphi(\bm{z}_{M})]$
@@
Covariance functions $\kappa_{\mathrm{pr}}(\bm{x},\bm{x}^{\prime})\coloneqq\langle\varphi(\bm{x}),\varphi(\bm{x}^{\prime})\rangle_{\mathbb{H}}$
@@
Covariance matrices $\mathbf{K}_{ZZ}\coloneqq\kappa_{\mathrm{pr}}(\mathbf{Z},\mathbf{Z})$
@@
Covariance operators $\Xi_{XX}\coloneqq\kappa_{\mathrm{pr}}(\mathbf{X},\mathbf{X})\otimes\mathcal{C}_{\textrm{bc}}$
@@
Feature vectors $\mathbf{k}_{Z}(\bm{x})\coloneqq\Phi_{Z}^{\top}\varphi(\bm{x})$
@@
Gramian matrix $\mathbf{C}_{XX}\coloneqq\mathbf{K}_{ZX}\,\mathbf{K}_{ZX}^{\top}$
Stiffness matrix $\mathbf{C}_{XY}\coloneqq\mathbf{K}_{ZX}\,\mathbf{K}_{ZY}^{\top}$
Regularization $\tilde{\mathbf{K}}_{XX}\coloneqq{\mathbf{K}}_{XX}+\sigma_{Y}^{2}\mathbf{I}_{N}$
@@
We start with a full Bayesian treatment of the embedded Perron–Frobenius operator (PFO) (Klus et al., 2020c), which we denote by $\mathcal{P}_{\varepsilon}:\mathbb{H}\rightarrow\mathbb{H}$. The embedded Perron–Frobenius operator, is the adjoint of the kernel Koopman operator. Unlike the PFO which propagates densities, the embedded PFO can be thought of as a push-forward map that acts on representations of probability densities. In short, the idea of feature maps is extended to the space of probability distributions by thinking of $\varphi(\bm{x})\,\forall\,\bm{x}\in\mathbb{X}$ as an injective representation of a density function. See Appendix C for a more detailed discussion of kernel mean embeddings.
Let us consider the ‘implicitly’ lifted training data set obtained from a dynamical system as $\left(\Phi_{X},\Phi_{Y}\right)=\{\left(\varphi\left(\bm{x}_{i}\right),\varphi\left(\bm{y}_{i}\right)\right)\}_{i=1}^{N}$, where the targets $\{\bm{y}_{i}\}_{i=1}^{N}$ are produced by the discrete-time flow map $\bm{F}:\mathbb{X}\rightarrow\mathbb{X}$, and $\mathbb{X}\subseteq\mathbb{R}^{D}$. We further assume only isotropic measurement noise:
@@
For the sake of clarity, allow a brief oversimplification and consider the D=1D\!=\!1 case. In this setting, the feature map $\varphi(\bm{x})\!\coloneqq\kappa_{\textrm{pr}}(\bm{x},\cdot)\,\in\,\mathbb{H}$ is defined by the GP prior over a one-dimensional flow map, i.e. a real- and scalar-valued kernel $\kappa_{\textrm{pr}}:\mathbb{X}\times\mathbb{X}\rightarrow\mathbb{R}$.
@@
Lifted Observation Model. We assume that a noisy lifted target $\varphi(\bm{y})\in\mathbb{H}$ can be approximated by the sum of the latent noise-free evaluation $\mu_{Y\mid\bm{x}}\coloneqq\mathcal{P}_{\varepsilon}\;\varphi(\bm{x})\in\mathbb{H}$ and zero mean Gaussian process noise:
@@
where $\nu_{i}(\bm{x})\sim\mathcal{GP}\left(0,\sigma_{Y}^{2}\kappa_{\textrm{bc}}(\bm{x},\bm{x}^{\prime})\right)$ models noise in the (infinite-dimensional) RKHS $\mathbb{H}$.
The purpose of the Bayesian-consistency kernel $\kappa_{\textrm{bc}}:\mathbb{X}\times\mathbb{X}\rightarrow\mathbb{R}$ is to relate the observation model in the state space to its counterpart in the RKHS. It models the correlations between the ‘components’ of $\varphi(\cdot)$ such that consistency is maintained with the flow map, i.e., ensuring that that $\bm{F}(\cdot)$ is an observable of the Koopman operator. The random variable $\nu_{i}\in\mathbb{H}$ is defined by the covariance operator $\mathcal{C}_{\textrm{bc}}:\mathbb{H}\rightarrow\mathbb{H}$, which in practice is required to be self-adjoint and invertible, and represents the statistical properties of the noisy full-state observable within the RKHS. For more details refer to Appendix E.
Similar to scalar-valued GP regression, our prior beliefs about the statistical properties of the operator $\mu_{Y\mid\cdot}:\mathbb{X}\rightarrow\mathbb{H}$ are represented by a GP, $\mu_{Y\mid\cdot}\sim\mathcal{GP}^{\infty}\left(\mathcal{M}_{\textrm{pr}},{\xi}_{\textrm{pr}}\right)$ [^1] The superscript attached to $\mathcal{GP}$ indicates the dimensionality of the random object associated with the Gaussian process.. However, here we are defining the prior covariance ${\xi}_{\textrm{pr}}\colon\mathbb{X}\times\mathbb{X}\to\mathbb{H}\,\otimes\,\mathbb{H}$ with an operator-valued kernel (Kadri et al., 2016). Then, without loss of generality, we choose the mean operator $\mathcal{M}_{\textrm{pr}}:\mathbb{X}\rightarrow\mathbb{H}$ such that it produces an identically zero function. It follows that the prior distribution over the latent operator evaluations $\mu_{Y\mid\mathbf{X}}\coloneqq\mathcal{P}_{\varepsilon}\;\Phi_{X}\in\mathbb{H}^{1\times N}$ is
@@
where $\Xi_{XX}$ is the prior covariance evaluated at $\left(\mathbf{X},\,\mathbf{X}\right)$ and hence an element of $(\mathbb{H}\,\otimes\,\mathbb{H})^{N\times N}$.
The Riesz-–Fréchet representation theorem (Hsing and Eubank, 2015; Fréchet, 1904) says that for each $\bm{x}\!\in\!\mathbb{X}$ and $g\!\in\!\mathbb{H}$, there is an element $\Xi(\bm{x})g\!:\mathbb{X}\rightarrow\mathbb{H}$, such that the reproducing property is
@@
This implies that the feature map $\Xi(\bm{x})$ is a bounded linear operator, $\Xi(\bm{x})\in\mathfrak{B}(\mathbb{H},\mathbb{H}^{\mathbb{X}})$, where $\mathbb{H}^{\mathbb{X}}$ denotes the $\mathbb{H}$-valued RKHS on $\mathbb{X}$ (i.e., $\mu_{Y\mid\cdot}\in\mathbb{H}^{\mathbb{X}}$). Thereby, the relationship between the reproducing kernel $\xi_{\textrm{pr}}(\cdot,\cdot)$ and its associated feature map $\Xi(\bm{x})$ is
@@
The entry at $\left(\xi_{\textrm{pr}}\left(\bm{x},\bm{x}^{\prime}\right)\right)_{\bm{y},\bm{y}^{\prime}}$ determines the prior covariance between $\mu_{Y\mid\bm{x}}(\bm{y})$ and $\mu_{Y\mid\bm{x}^{\prime}}(\bm{y}^{\prime})$, i.e., it expresses the degree to which the response at $(\bm{x},\bm{y})$ is affected by $(\bm{x}^{\prime},\bm{y}^{\prime})$ — before incorporating data.
Definition 2.1 models the lifted targets as scalar-valued GPs, and therefore the likelihood can be expressed as $p(\varphi(\bm{y})\mid\mathcal{P}_{\varepsilon},\bm{x})=\mathcal{GP}(\mu_{Y\mid\bm{x}},\sigma_{Y}^{2}\kappa_{\textrm{bc}})$. Thereby, the posterior retains the same structure as in the scalar-valued case (Alvarez et al., 2012; Micchelli and Pontil, 2004, 2005):
$\mathcal{M}_{\textrm{pst}}(\bm{x})$ $\coloneqq\mathcal{A}\;\xi_{X}({\bm{x}})=\sum_{i=1}^{N}\mathcal{A}_{i}\;{\xi}_{\textrm{pr}}(\bm{x}_{i},\bm{x}),$ (2.1)
@@
We unpack these equations by concatenating the noisy targets ‘end-to-end’ in $\Upsilon^{\top}\!\coloneqq\!\operatorname{vec}(\Phi_{Y})\!\in\!\mathbb{H}^{\oplus N}$. We denote the weight operator by $\mathcal{A}\coloneqq\Upsilon\;\tilde{\Xi}_{XX}^{-1}$, the noisy covariance operator by $\tilde{\Xi}_{XX}\coloneqq\Xi_{XX}+\Sigma$, and the feature “vector" by $\mathbf{\xi}_{X}(\bm{x})\coloneqq[{\xi}_{\textrm{pr}}(\bm{x}_{1},\bm{x}),\ldots,{\xi}_{\textrm{pr}}(\bm{x}_{N},\bm{x})]^{\top}$.
We will elaborate on this point in Section 2.4, but the structure the operator-valued kernel is restricted to $\xi_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\!=\!\kappa_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\,\mathcal{C}_{\textrm{bc}}$. This form allows us to express the covariance operators in terms of the prior covariances over latent values of the flow map and Kronecker tensor products with $\mathcal{C}_{\textrm{bc}}$, i.e., $\Xi_{XX}\coloneqq\mathbf{K}_{XX}\otimes\mathcal{C}_{\textrm{bc}}$ where $\mathbf{K}_{XX}\coloneqq\Phi_{X}^{\top}\,\Phi_{X}\in\mathbb{R}^{N\times N}$. Similarly, the “sensor" noise covariance operator is denoted by $\Sigma\coloneqq\mathbf{I}_{N}\,\otimes\,\sigma_{Y}^{2}\,\mathcal{C}_{\textrm{bc}}$, and the feature vector is $\mathbf{\xi}_{X}(\bm{x})=\mathbf{k}_{X}(\bm{x})\otimes\mathcal{C}_{\textrm{bc}}$ where $\mathbf{k}_{X}(\bm{x})\coloneqq\Phi_{X}^{\top}\varphi(\bm{x})$. Thereby, in essence we are implementing an intrinsic model of coregionalization (Alvarez et al., 2012).
@@
Hence, we recover the same expression for ${\mathcal{P}}_{\varepsilon}$ derived in (Klus et al., 2020c), while also mirroring the frequentist approach of (Grünewälder et al., 2012). In the expression of the mean, the operator $\mathcal{C}_{\textrm{bc}}$ plays no role whatsoever, yet we observe that the posterior covariance over ${\mathcal{P}}_{\varepsilon}\varphi(\bm{x})$ retains the same form as the prior: $\xi_{\textrm{pst}}(\bm{x},\bm{x}^{\prime})=\kappa_{\textrm{pst}}(\bm{x},\bm{x}^{\prime})\,\mathcal{C}_{\textrm{bc}}$, i.e., a linear transformation of the covariance over the flow map.
In the case where $D>1$, we may want to treat the prior and Bayesian-consistency kernels as matrix-valued and include anisotropic measurement noise (Alvarez et al., 2012; Wilson et al., 2012). However, for the rest of this manuscript we will continue assuming that the channels in the flow map are unrelated and that the sensor noise is isotropic.